OpenStudy (anonymous):

Red and gray bricks were used to build a decorative wall. The number of red bricks/the number of gray bricks is 5/2. There were 210 bricks used in all. How many red bricks were used? answer choices: a.150 b.60 c.42 d.30

OpenStudy (anonymous):

@TheRealMeeeee @xo_kansasprincess_xo @abbles @elementwielder @iambatman

OpenStudy (therealmeeeee):

I'm thinkiing A but I might be wrong

OpenStudy (anonymous):

ok thanks anyway

OpenStudy (ranga):

Number of red bricks / Number of gray bricks = 5/2 That means if there are 7 bricks total, 5 will be red and 2 will be gray. So if the total number of bricks is 7, the number of red bricks is 5. If the total number of bricks is 210, the number of red bricks will be ? (use proportions to solve).

OpenStudy (anonymous):

the answer is A just put the answers in divided by what is left out of the total and keep it in fraction form when you get an answer

OpenStudy (therealmeeeee):

It's A

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Let r = red and g = gray. r/g = 5/2 or r = 5g/2. It tells you that r + g = 210. Substitute the value for r (r = 5g/2) into the equation r + g = 210, which gives you: 5g/2 + g = 210. Solve this equation for g. Once you know g, you can find r.

OpenStudy (anonymous):

@mimipink16 didnt you ask this question before?